OpenStudy (anonymous):

Given g(x) = x+1 x−1 , solve g(x) = 0

OpenStudy (saifoo.khan):

Since g(x) = 0 g(x) = x+1 0 = x+ 1 Solve.

jhonyy9 (jhonyy9):

can you rewrite it ? what is x-1 ?

OpenStudy (anonymous):

Sorry given g(x)= x+1/x-1, solve g(x)=0

OpenStudy (saifoo.khan):

\[g(x) = \frac{x+1}{x-1}\] \[0 = \frac{x+1}{x-1}\]

jhonyy9 (jhonyy9):

ok than first of all you need push the restriction for x when the denominator will be equal zero

jhonyy9 (jhonyy9):

because than the fraction will be undefined

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

so what will be ?

jhonyy9 (jhonyy9):

Rose ! are you there ?